# Revisiting the hydrodynamic modulation of short surface waves by longer waves

### v2

@@ Abstract
Hydrodynamic modulation of short ocean surface waves by longer ambient waves significantly influences remote sensing, interpretation of in situ wave measurements, and numerical wave forecasting. This paper revisits the wave crest and action conservation laws and derives steady, nonlinear, analytical solutions for the change of short-wave wavenumber, action, and gravitational acceleration due to the presence of longer waves. We validate the analytical solutions with numerical simulations of the full conservation equations. The nonlinear analytical solutions of short-wave wavenumber, amplitude, and steepness modulation significantly deviate from the linear analytical solutions of Longuet-Higgins & Stewart (1960), and are similar to the nonlinear numerical solutions by Longuet-Higgins (1987) and Zhang & Melville (1990). The short-wave steepness modulation is attributed 5/8 to wavenumber, 1/4 due to wave action, and 1/8 due to effective gravity. Examining the homogeneity and stationarity requirements for the conservation of wave action reveals that stationarity is a stronger requirement and is generally not satisfied for very steep long waves. We examine the results of Peureux et al. (2021) who found through numerical simulations that the short-wave modulation grows unsteadily with each long-wave passage. We show that this unsteady growth only occurs for homogeneous initial conditions as a special case and not generally. The proposed steady solutions are a good approximation of the nonlinear crest-action conservation solutions in long-wave steepness $\lesssim 0.2$. Except for a subset of initial conditions, the solutions to the non-linearised crest-action conservation equations are mostly steady in the reference frame of the long waves.
@@ 2 Governing equations
where $\omega$ is the angular frequency in a fixed reference frame, $g$ is the gravitational acceleration, $k$ is the wavenumber, and $U$ is the mean advective current in the direction of the wave propagation. A current in the direction of the waves increases their absolute frequency without changing their wavenumber. An important limitation here is that $U$ must be slowly varying on the scales of the wave period (Bretherton & Garrett, 1968). From the perspective of short waves riding on long waves, the advective current is the horizontal near-surface orbital velocity of the long wave. Strictly, a velocity profile must be considered when determining the advective velocity for a wave of any given wavenumber (Stewart & Joy, 1974), however, for brevity, here we will assume a surface velocity to be the advective one for short waves. The evolution of the short-wave wavenumber is described by the conservation of wave crests (e.g., Whitham, 1974; Phillips, 1981):
@@
where $N$ is the action of short waves defined as the ratio of their energy to their intrinsic frequency, $E/\sigma$, and $C_{g}=1/2\sqrt{g/k}$ is their group speed in deep water. The wave action expressed as $E/\sigma$ is commonly taken as conservative in wave modulation theory (Phillips, 1981; Longuet-Higgins, 1987), however as Dysthe et al. (1988) show, it is conservative only for gently sloped waves. We describe this and other requirements in more detail further below and show in what long-wave conditions they are satisfied. The wave crest conservation (3) follows from the assumption of a conserved wave phase, which is limited by the same set of requirements as those for the wave action conservation (e.g., Whitham, 1974). Although the absence of wave growth and dissipation is not generally realistic in the open ocean, it is a useful approximation here to isolate the effects of hydrodynamic modulation solely due to the motion of the long waves.
@@
This equation is similar to (6), except for the last term which represents the inhomogeneity of the group speed. As Peureux et al. (2021) explained, the presence of this term causes unsteady growth of wave action in infinite long-wave trains. (6) and (7) are the governing equations that we approximate and solve analytically in §3, and numerically in their full form in §4. This system of equations is semi-coupled, meaning that it is possible to solve (6) on its own, but solving (7) requires also the solution of (6). In other words, the wave kinematics (the wavenumber distribution) are not concerned with the wave action. In contrast, the wave action is governed, among others, by the group velocity and thus depends on the wavenumber.
@@
Evaluating these velocities at the wave surface $z=\eta_{L}$ rather than at the mean water level $z=0$ is a departure from the linear theory, making the treatment of the long waves here as weakly nonlinear. As Zhang & Melville (1990) point out, this is appropriate to do when the amplitude of the long waves is of the similar magnitude or larger than the wavelength of the short waves. However, evaluating these velocities at the free surface $z=\eta_{L}$ introduces an additional error of $\mathcal{O}(\varepsilon_{L}^{2})$. The surface velocity errors are described in more detail in Appendix A.
@@ 3.2 Gravity modulation
Let us now examine the contributions of the different terms to the short-wave effective gravity modulation. Fig. 3a shows the effective gravity modulation of short waves as a function of the long-wave phase for $\varepsilon_{L}=0.3$, for a variety of the gravity modulation forms. Although quite steep, this value of $\varepsilon_{L}$ allows easier visualisation of the modulation differences. To $\mathcal{O}(\varepsilon_{L})$, the gravity modulation is out of phase with the long-wave elevation, with the short waves experiencing a lower effective gravity on the long-wave crests and a higher effective gravity in the troughs. Evaluating the orbital accelerations at $z=\eta$ and the Lagrangian gravity correction each have an $\mathcal{O}(\varepsilon_{L}^{2})$ contribution but of opposite signs. The former amplifies the gravity modulation while the latter attenuates it. The curvilinear effects are significant at the front and rear faces of the long wave but not near the crests and troughs. The long-wave surface slope has a $\mathcal{O}(\varepsilon_{L}^{2})$ effect. Naturally, the slope is exactly zero at the crests and the troughs, and its effects are largest at the front and rear faces of the long wave, where the covariance between the local slope and the horizontal accelerations is largest. Evaluating the velocities and accelerations at the surface of a third-order Stokes wave causes a correction mostly at the faces of the long wave. Finally, numerically computing the surface velocities and accelerations of a fully-nonlinear wave allows us to exclude uncertainties associated with the truncation errors of the Stokes expansion at first and third orders in $\varepsilon_{L}$. At this steepness, the fully-nonlinear solution is marginally different from that of the third-order Stokes wave (see also Appendix A for the errors in surface velocity estimates of the linear and third-order Stokes approximations).
@@ 3.4 Limits to the wave action conservation
Bretherton & Garrett (1968) also require that the medium is nearly uniform over the vertical scales of motion of the short waves. For the ambient velocity being the orbital velocity of the long wave, this is trivially satisfied with (33).
Figure 6: Homogeneity (left) and stationarity (right) of the short-wave wavenumber (top) and effective gravity (bottom) as a function of the long-wave steepness $\varepsilon_{L}$ and the wavenumber ratio $k/k_{L}$, based on the linearised solutions (37-40). Solid and dashed lines highlight the 0.99 and 0.90 values, respectively.
@@
Fig. 6 shows the homogeneities and stationarities of the short-wave wavenumber and effective gravity based on the steady solutions and their derived criteria as minima of (29)-(30). Stationarity is a stronger requirement than homogeneity. At the lowest scale separation considered ($k/k_{L}=10$), short-wave wavenumber, action, and effective gravity are all strongly homogeneous ($H_{k,N,g}>0.99$) for $\varepsilon_{L}<0.1$. However, the solutions are only strongly stationary ($S_{k,N,g}>0.99$) at a high scale separation of $k/k_{L}=100$ for $\varepsilon_{L}<0.1$. If we consider weakly stationary conditions as $S_{k,N,g}>0.9$, then the solutions satisfy the requirements for the wave action conservation for $\varepsilon_{L}<0.3$ at $k/k_{L}=10$, and for $\varepsilon_{L}<0.4$ at $k/k_{L}>20$. Although Bretherton & Garrett (1968) state that the wave action balance is valid to $\mathcal{O}(k_{L}/k)$, applying the general criteria (29)-(30) yields an error of $\mathcal{O}(\varepsilon_{L}k_{L}/k)$. We will revisit these criteria based on the numerical solutions of nonlinear wave action and crest conservation equations in the next section.
@@ 6 Conclusions
In this paper we revisited the problem of short-wave modulation by long waves from a wave action conservation perspective. Using the linearised wave crest and action conservation equations coupled with nonlinear solutions for the effective gravity, we derived steady, nonlinear, analytical solutions for the modulation of short wave wavenumber, amplitude, steepness, intrinsic frequency, and phase speed. The latter two were previously examined in the context of short-wave modulation (Longuet-Higgins, 1962; Longuet-Higgins & Phillips, 1962), and may be important for correctly calculating the wind input into short waves, which is proportional to the wind speed relative to the wave celerity. The approximate steady solutions yield higher wavenumber and amplitude modulation than that of Longuet-Higgins & Stewart (1960), and somewhat lower than the numerical solutions of Longuet-Higgins (1987) and Zhang & Melville (1990). This is mainly due to the analytical solution requiring the crest and action conservation equations to be linearised, and the nonlinearity arising due to the evaluation of short waves on the surface of the long wave (as opposed to the mean water level) and due to nonlinear effective gravity. In the steady solutions, the short-wave steepness modulation is dominated by the wavenumber modulation (5/8), followed by the wave action modulation (1/4), and is least affected by the effective gravity modulation (1/8). The homogeneity and stationarity criteria for wave action conservation are evaluated for both the analytical and numerical solutions. The stationarity is a stronger criterion than homogeneity, and is found to be strongly satisfied ($>99\%$ stationary) for very gently sloped waves and weakly satisfied ($>90\%$ stationary) for steeper waves.
@@
Appendix A Surface velocity errors

### v3

@@ Abstract
Hydrodynamic modulation of short ocean surface waves by longer ambient waves significantly influences remote sensing, interpretation of in situ wave measurements, and numerical wave forecasting. This paper revisits the wave crest and action conservation laws and derives steady, nonlinear, analytical solutions for the change of short-wave wavenumber, action, and gravitational acceleration due to the presence of longer waves. We validate the analytical solutions with numerical solutions of the full crest and action conservation equations. The nonlinear analytical solutions of short-wave wavenumber, amplitude, and steepness modulation significantly deviate from the linear analytical solutions of Longuet-Higgins & Stewart (1960), and are similar to the nonlinear numerical solutions by Longuet-Higgins (1987) and Zhang & Melville (1990). The short-wave steepness modulation is attributed 5/8 to wavenumber, 1/4 due to wave action, and 1/8 due to effective gravity. Examining the homogeneity and stationarity requirements for the conservation of wave action reveals that stationarity is a stronger requirement and is generally not satisfied for very steep long waves. We examine the results of Peureux et al. (2021) who found through numerical simulations that the short-wave modulation grows unsteadily with each long-wave passage. We show that this unsteady growth only occurs for homogeneous initial conditions as a special case and not generally. The proposed steady solutions are a good approximation of the nonlinear crest-action conservation solutions in long-wave steepness $\lesssim 0.2$. Except for a subset of initial conditions, the solutions to the non-linearised crest-action conservation equations are mostly steady in the reference frame of the long waves.
@@ 2 Governing equations
where $\omega$ is the angular frequency in a fixed reference frame, $g$ is the gravitational acceleration, $k$ is the wavenumber, and $U$ is the mean advective current in the direction of the wave propagation. A current in the direction of the waves increases their absolute frequency without changing their wavenumber. An important limitation here is that $U$ must be slowly varying on the scales of the wave period (Bretherton & Garrett, 1968). From the perspective of short waves riding on long waves, the advective current is the horizontal near-surface orbital velocity of the long wave. Strictly, shear must be considered when determining the advective velocity for a wave of any given wavenumber (Stewart & Joy, 1974; Ellingsen, 2014). Here, we assume the surface velocity to be the advective one for short waves, and demonstrate later in §$\S$3.4 that this is valid for sufficient scale separation between the short and long waves. The evolution of the short-wave wavenumber is described by the conservation of wave crests (e.g., Whitham, 1974a; Phillips, 1981):
@@
where $N$ is the action of short waves defined as the ratio of their energy to their intrinsic frequency, $E/\sigma$, and $C_{g}=1/2\sqrt{g/k}$ is their group speed in deep water. The wave action expressed as $E/\sigma$ is commonly taken as conservative in wave modulation theory (Phillips, 1981; Longuet-Higgins, 1987), however as Dysthe et al. (1988) show, it is conservative only for gently sloped waves. We describe this and other requirements in more detail further below and show in what long-wave conditions they are satisfied. The wave crest conservation (3) follows from the assumption of a conserved wave phase, and is another prerequisite for wave action conservation (e.g., Whitham, 1974a). Although the absence of wave growth and dissipation is not generally realistic in the open ocean, it is a useful approximation here to isolate the effects of hydrodynamic modulation solely due to the motion of the long waves. The governing equations (2)-(4) are derived in Appendix A.
@@
This equation is similar to (6), except for the last term which represents the inhomogeneity of the group speed. As Peureux et al. (2021) explained, the presence of this term causes unsteady growth of wave action in infinite long-wave trains. (6) and (7) are the governing equations that we approximate and solve analytically in §3, and numerically in their full form in §4. This system of equations is semi-coupled, meaning that it is possible to solve (6) on its own, but solving (7) requires also the solution of (6). In other words, the wave kinematics (the wavenumber distribution) are not concerned with the wave action. In contrast, the wave action is governed, among others, by the group velocity and thus depends on the wavenumber. This convenient property is only possible if we consider the linear form of the dispersion relation, that is, if the series are truncated at the second-order solution in the Stokes expansion. For a third-order expansion, an additional term that is proportional to the square of the wave steepness appears in the dispersion relation. In that case, equations (6) and (7) become fully coupled and the analytical solution for the wavenumber is no longer feasible.
@@
Evaluating these velocities at the wave surface $z=\eta_{L}$ rather than at the mean water level $z=0$ is a departure from the linear theory, making the treatment of the long waves here as weakly nonlinear. As Zhang & Melville (1990) point out, this is appropriate to do when the amplitude of the long waves is of the similar magnitude or larger than the wavelength of the short waves. However, evaluating these velocities at the free surface $z=\eta_{L}$ introduces an additional error of $\mathcal{O}(\varepsilon_{L}^{2})$. The surface velocity errors are described in more detail in Appendix B.
@@ 3.2 Gravity modulation
Let us now examine the contributions of the different terms to the short-wave effective gravity modulation. Fig. 3a shows the effective gravity modulation of short waves as a function of the long-wave phase for $\varepsilon_{L}=0.3$, for a variety of the gravity modulation forms. Although quite steep, this value of $\varepsilon_{L}$ allows easier visualisation of the modulation differences. To $\mathcal{O}(\varepsilon_{L})$, the gravity modulation is out of phase with the long-wave elevation, with the short waves experiencing a lower effective gravity on the long-wave crests and a higher effective gravity in the troughs. Evaluating the orbital accelerations at $z=\eta$ and the Lagrangian gravity correction each have an $\mathcal{O}(\varepsilon_{L}^{2})$ contribution but of opposite signs. The former amplifies the gravity modulation while the latter attenuates it. The curvilinear effects are significant at the front and rear faces of the long wave but not near the crests and troughs. The long-wave surface slope has a $\mathcal{O}(\varepsilon_{L}^{2})$ effect. Naturally, the slope is exactly zero at the crests and the troughs, and its effects are largest at the front and rear faces of the long wave, where the covariance between the local slope and the horizontal accelerations is largest. Evaluating the velocities and accelerations at the surface of a third-order Stokes wave causes a correction mostly at the faces of the long wave. Finally, numerically computing the surface velocities and accelerations of a fully-nonlinear wave allows us to exclude uncertainties associated with the truncation errors of the Stokes expansion at first and third orders in $\varepsilon_{L}$. At this steepness, the fully-nonlinear solution is marginally different from that of the third-order Stokes wave (see also Appendix B for the errors in surface velocity estimates of the linear and third-order Stokes approximations).
@@ 3.4 Limits to the wave action conservation
Bretherton & Garrett (1968) also require that the medium is nearly uniform over the vertical scales of motion of the short waves:
$\frac{1}{U}\frac{\partial U}{\partial z}\ll\widetilde{k}$ (35)
which, for a long deep water wave, is equivalent to (33). Alternatively, if we use the second Froude number criterion following Ellingsen (2014) and quantify the shear length under the long wave surface as:
$l_{S}=\frac{g}{S^{2}}\approx\frac{g}{\varepsilon_{L}^{2}\omega_{L}^{2}}$ (36)
we arrive to an additional requirement:
$\frac{k}{k_{L}}\gg\varepsilon_{L}^{2}$ (37)
which relates the scale separation to the long-wave steepness.
Figure 6: Homogeneity (left) and stationarity (right) of the short-wave wavenumber (top) and effective gravity (bottom) as a function of the long-wave steepness $\varepsilon_{L}$ and the wavenumber ratio $k/k_{L}$, based on the linearised solutions (40-43). Solid and dashed lines highlight the 0.99 and 0.90 values, respectively.
@@
Fig. 6 shows the homogeneities and stationarities of the short-wave wavenumber and effective gravity based on the steady solutions and their derived criteria as minima of (29)-(30). Stationarity is a stronger requirement than homogeneity. At the lowest scale separation considered ($k/k_{L}=10$), short-wave wavenumber, action, and effective gravity are all strongly homogeneous ($H_{k,N,g}>0.99$) for $\varepsilon_{L}<0.1$. However, the solutions are only strongly stationary ($S_{k,N,g}>0.99$) at a high scale separation of $k/k_{L}=100$ for $\varepsilon_{L}<0.1$. If we consider weakly stationary conditions as $S_{k,N,g}>0.9$, then the solutions satisfy the requirements for the wave action conservation for $\varepsilon_{L}<0.3$ at $k/k_{L}=10$, and for $\varepsilon_{L}<0.4$ at $k/k_{L}>20$. Although Bretherton & Garrett (1968) state that the wave action balance is valid to $\mathcal{O}(k_{L}/k)$, applying the general criteria (29)-(30) yields an error of $\mathcal{O}(\varepsilon_{L}k_{L}/k)$. We will revisit these criteria based on the numerical solutions of nonlinear wave action and crest conservation equations in the next section. For the asymptotic limits of wave action conservation in terms of the wave short-wave steepness, see Appendix A.3.
@@ 6 Conclusions
In this paper we revisited the problem of short-wave modulation by long waves from a wave action conservation perspective. Using the linearised wave crest and action conservation equations coupled with nonlinear solutions for the effective gravity, we derived steady, nonlinear, analytical solutions for the modulation of short wave wavenumber, amplitude, steepness, intrinsic frequency, and phase speed. The latter two were previously examined in the context of resonant interactions (Longuet-Higgins, 1962; Longuet-Higgins & Phillips, 1962), a modulation process distinct from the one discussed here. Both effects, however, may be important for correctly calculating the wind input into short waves, which is proportional to the wind speed relative to the wave celerity. The approximate steady solutions yield higher wavenumber and amplitude modulation than that of Longuet-Higgins & Stewart (1960), and somewhat lower than the numerical solutions of Longuet-Higgins (1987) and Zhang & Melville (1990). This is mainly due to the analytical solution requiring the crest and action conservation equations to be linearised, and the nonlinearity arising due to the evaluation of short waves on the surface of the long wave (as opposed to the mean water level) and due to nonlinear effective gravity. In the steady solutions, the short-wave steepness modulation is dominated by the wavenumber modulation (5/8), followed by the wave action modulation (1/4), and is least affected by the effective gravity modulation (1/8). The homogeneity and stationarity criteria for wave action conservation are evaluated for both the analytical and numerical solutions. The stationarity is a stronger criterion than homogeneity, and is found to be strongly satisfied ($>99\%$ stationary) for very gently sloped waves and weakly satisfied ($>90\%$ stationary) for steeper waves.
@@
Appendix A Derivation of the governing equations
Here we derive the governing equations for the short waves (2)-(4) and discuss their asymptotic consistency.
A.1 Wave dispersion
An irrotational flow has a velocity potential such that the continuity can be expressed as:
$\nabla^{2}\phi=0$ (44)
Boundary conditions that allow solving for $\phi$ are the dynamic and kinematic free surface boundary conditions, respectively:
$\frac{\partial\phi}{\partial t}+\frac{1}{2}(\nabla{\phi})^{2}+gz=0\quad\text{at}\quad z=\eta(x,t)$ (45)
$\frac{\partial\eta}{\partial t}+\frac{\partial\phi}{\partial x}\frac{\partial\eta}{\partial x}+\frac{\partial\phi}{\partial z}=0\quad\text{at}\quad z=\eta(x,t)$ (46)
Also required is the bottom boundary condition:
$\frac{\partial\phi}{\partial z}=0\quad\text{at}\quad z=-h$ (47)
In the Stokes (1847) perturbation approach, the velocity potential and elevation are expanded in a power series of the wave steepness $\varepsilon=ak$ as:
$\phi=\sum_{n=1}^{\infty}\varepsilon^{n}\phi_{n}$ (48)
$\eta=\sum_{n=1}^{\infty}\varepsilon^{n}\eta_{n}$ (49)
The power series expansion requires $\varepsilon\ll 1$. The velocity potential and elevation series are truncated at the desired order and inserted into the dynamic free surface boundary condition (45). Up to the third order, the solutions for the velocity potential and the corresponding velocity components are (8)-(10), and the surface elevation is (21). Evaluating these quantities to higher orders in $\varepsilon$ is possible at the expense of higher complexity and with diminishing returns in accuracy.
Combining the solutions for the potential and the surface elevation with the kinematic free surface boundary condition (46) yields the dispersion relation, which in the linear case and in deep water is (2). At the third order in $\varepsilon$, the frequency begins to depend on the wave steepness $\varepsilon$ as well:
$\omega=\sqrt{gk}\left(1+\frac{1}{2}\varepsilon^{2}+\mathcal{O}(\varepsilon^{4})\right)+Uk$ (50)
[…]
A.2 Conservation of wave crests
A system with a well defined phase function $\psi(x,t)$ also has the wavenumber and frequency defined as its gradients with respect to space and time, respectively:
$k=\frac{\partial\psi}{\partial x}$ (51)
$\omega=-\frac{\partial\psi}{\partial t}$ (52)
implying $\psi=kx-\omega t$. Differentiating (51)-(52) with respect to time and space, respectively, and then eliminating $\psi$ yields:
$\frac{\partial k}{\partial t}+\frac{\partial\omega}{\partial x}=0$ (53)
which is the crest conservation equation (3). That this is the wavenumber conservation equation with the group speed as the transport velocity can be recognized by writing:
$\frac{\partial k}{\partial t}+\frac{\partial\omega}{\partial k}\frac{\partial k}{\partial x}=0$ (54)
This conservation equation is valid regardless of the order of nonlinearity considered, as long as $\psi$ is well defined. Nonlinearity in terms of $\varepsilon$ can, however, enter (53) through nonlinear forms of $\omega$ such as (50).
A.3 Conservation of wave action
The conservation of wave action (4) is most readily derived following the variational approach by Whitham (1965). Begin with a statement that infinitesimal variations of a Lagrangian of a system over a spatial domain and time interval must vanish:
[…]
Here, $L$ is the instantaneous Lagrangian density, a scalar quantity that corresponds to the difference between the kinetic and potential energy of the system:
[…]
Luke (1967) showed that the equations for irrotational surface waves follow from the variational approach if the Lagrangian $L$ is defined instead as:
[…]
(56) and (57) are equivalent in the sense that they both yield (44), however, varying the latter is necessary to attain the correct boundary conditions (Whitham, 1967). Notice that for water waves the instantaneous Lagrangian being zero is equivalent to the Bernoulli’s equation, which is also the dynamic free surface boundary condition.
Then, Whitham’s averaged variation principle is obtained by averaging the instantaneous Lagrangian $L$ over one wave period:
[…]
where $\mathcal{L}$ is the averaged Lagrangian density. The averaged variational principle then states:
[…]
The averaged Lagrangian for water waves is obtained by integrating (57) over depth, and then averaging over the phase (58). In the deep water limit, it is:
$\mathcal{L}=\frac{1}{4}\left(1-\frac{\omega^{2}}{gk}\right)ga^{2}+\mathcal{O}(\varepsilon^{4})$ (60)
[…]
$\mathcal{L}=\frac{1}{2}\left(1-\frac{\omega^{2}}{gk}\right)E+\mathcal{O}(\varepsilon^{4})$ (61)
Whitham’s conservation equations for water waves are:
$\frac{\partial\mathcal{L}}{\partial a}=0$ (62)
$\frac{\partial}{\partial t}\left(\frac{\partial\mathcal{L}}{\partial\omega}\right)-\frac{\partial}{\partial x}\left(\frac{\partial\mathcal{L}}{\partial k}\right)=0$ (63)
Inserting (61) into (63) and defining:
$\frac{\partial\mathcal{L}}{\partial\omega}=\frac{E}{\sigma}\equiv N$ (64)
yields the conservation of wave action:
$\frac{\partial}{\partial t}\left[N+\mathcal{O}(\varepsilon^{4})\right]+\frac{\partial}{\partial x}\left[(C_{g}+\mathcal{O}(\varepsilon^{2}))N+\mathcal{O}(\varepsilon^{4})\right]=0$ (65)
Due to the $\mathcal{O}(\varepsilon^{2})$ correction to the group speed, the overall truncation error for the wave action balance is $\mathcal{O}(\varepsilon^{2})$ as well:
$\frac{\partial N}{\partial t}+\frac{\partial(C_{g}N)}{\partial x}=\mathcal{O}(\varepsilon^{2})$ (66)
which was introduced as a governing equation in §2. Although the truncation error of $\mathcal{O}(\varepsilon^{2})$ arising from $C_{g}$ seems to have been omitted in the original derivation by Whitham (1967), it appeared correctly later in Whitham (1974b). An important nuance about applying the nonlinear dispersion here is that the secondary (wave-induced) mean flow and set-up effects can be ignored in the deep water limit; this is discussed by Whitham (1974a). The terms proportional to $\varepsilon^{4}$ do not appear in the averaged Lagrangian (61) until the third-order Stokes waves are considered, for which the dispersion gains a $\mathcal{O}(\varepsilon^{2})$ correction. In that case, the conserved quantity is:
$\frac{\partial\mathcal{L}}{\partial\omega}=\frac{E}{\sigma}(1+\frac{1}{2}\varepsilon^{2})$ (67)
and the group speed gains a nonlinear factor in $\varepsilon^{2}$ as well. (66), then, maintains the same form, but with the truncation error of $\mathcal{O}(\varepsilon^{4})$. This derivation, as well as the averaged Lagrangian principle applied to the wave-induced mean flow over arbitrary water depths, is detailed at greater length by Whitham (1967).
To consider the effect of slowly varying ambient currents, Bretherton & Garrett (1968) carried out the same derivation except that they propagated the ambient current derivatives $\partial U/\partial t$ and $\partial U/\partial x$ through the averaged Lagrangian. To allow these derivatives to drop out of the averaging integral, the authors make an argument that the variation must be sufficiently small, no larger than $\mathcal{O}(\frac{1}{\sigma U}\frac{\partial U}{\partial t})$ and $\mathcal{O}(\frac{1}{kU}\frac{\partial U}{\partial x})$, respectively. Taking into consideration the asymptotic limits in terms of the wave steepness $\varepsilon$ as well as the variation of the ambient current, the wave action balance can be written as:
$\frac{\partial N}{\partial t}+\frac{\partial}{\partial x}\left[\left(C_{g}+U\right)N\right]=\mathcal{O}\left(\varepsilon^{4},\frac{1}{\sigma U}\frac{\partial U}{\partial t},\frac{1}{kU}\frac{\partial U}{\partial x}\right)$ (68)
§3.4 deals specifically with quantifying the ambient velocity variation in the context of long-wave orbital velocities.
There is an alternative derivation of the wave action conservation that was pointed out by Bretherton & Garrett (1968). Starting from the wave energy conservation in non-uniform currents derived by Longuet-Higgins & Stewart (1961)
$\frac{\partial E}{\partial t}+\frac{\partial}{\partial x}\left[\left(C_{g}+U\right)E\right]+S_{x}\frac{\partial U}{\partial x}=0$ (69)
where $S_{x}$ is the wave radiation stress, and combining it with the dispersion relation (2) and conservation of crests (3), they show that (69) is equivalent to the wave action conservation (4). This derivation is found in the Appendix of Bretherton & Garrett (1968), and was later recognized by Whitham (1974b) as well.
In summary, asymptotically consistent sets of dispersion and action conservation equations vary depending on the level of nonlinearity considered. In the case of linear waves, the dispersion relation (2), the averaged Lagrangian (61), the adiabatic quantity (64) and the wave action conservation (66) are asymptotically consistent. For third-order Stokes waves, the dispersion relation (50), the adiabatic quantity (67), together with the group speed that includes a $\mathcal{O}(\varepsilon^{2})$ correction, are governed by the action conservation law (66) and are asymptotically consistent. This is summarized in Table 2.
Order Frequency Group speed Wave action conservation
1, 2 $\sqrt{gk}(1+\mathcal{O}(\varepsilon^{2}))$ $\frac{1}{2}\sqrt{\frac{g}{k}}(1+\mathcal{O}(\varepsilon^{2}))$ $\frac{\partial N}{\partial t}+\frac{\partial(C_{g}N)}{\partial x}=\mathcal{O}(\varepsilon^{2})$
3 $\sqrt{gk}(1+\frac{1}{2}\varepsilon^{2}+\mathcal{O}(\varepsilon^{4}))$ $\frac{1}{2}\sqrt{\frac{g}{k}}(1+\frac{5}{2}\varepsilon^{2}+\mathcal{O}(\varepsilon^{4}))$ $\frac{\partial N}{\partial t}+\frac{\partial(C_{g}N)}{\partial x}=\mathcal{O}(\varepsilon^{4})$
Table 2: Asymptotically consistent sets of dispersion and wave action relations for deep water and their respective truncation errors in terms of the wave steepness $\varepsilon$, for the first three Stokes expansion orders.
Appendix B Surface velocity errors